How many years will it take $10,000 to grow into one million dollars if it is invested at 4.35% compounded semi-annually?

  • A. 216.31 years

  • B. 107.01 years

  • C. 214.03 years

  • D. 108.15 years
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,000,000=10,000*(1+0.0435/2)^2n

(1,000,000/10,000)=1.02175^2n

Taking log on both sides;

log 100=2n*log 1.02175

n=1/2[log 100/log 1.02175]

=107.01 years(Approx).